IlIlIcase ill alcuuis leolvavic (U, JUU Decrease in inventory Decrease in accounts payable 1,700 (3,600) $ 24,600 Cash flowsAvalos Corporation is preparing its annual financial statements at December 31 of the current year. Listed here are the itemsNeed help finding the net cash inflows for the year and beginning cash balance
Answer #1

Solution:

Net cash flow for the year = 24600 - 45900 + 62000 = $40,700

Beginning cash balance (Given in Question) = $30,000

Ending cash balance = $40700 + 30000 = $70,700